the max range of a particle with certain Speed on a horizontal plane is R. find the max range when projected on an incline with inclination of 30°

Grade:11

1 Answers

Eshan
5 years ago
Dear student,

Range of projectile is given asR=\dfrac{2u^2sin\theta cos\theta}{g}

Range is maximum for\theta=45^{\circ}

Therefore,\dfrac{R'}{R}=\dfrac{sin30^{\circ}cos30^{\circ}}{sin45^{\circ}cos45^{\circ}}=\dfrac{\sqrt{3}}{2}
\implies R'=\dfrac{\sqrt{3}R}{2}